Use euclide division algorithm to find hcf of 56 and 72 and hence express it in the form of 56x +72y

Grade:6

1 Answers

Shambhavi Mishra
6 years ago
First divide 72 by 56 we get the following algorithm
72=56*1+16
56=16*3+8
16=8*2+0
Since we got the hcf i.e 8 in the second step
Therefore by using the second step we get is
56-(16*3)=8
Putting value of 16 from step 1 in the above equation we get
56-((72-56)*3)=8
Opening brackets
56*4-3*72
56*4+72(-3)
Therefore
X=4 & y=(-3)
 
Check the soon
224+(-216)=8
 
Hence verified